Moller Polarimeter

From PolWiki
Jump to: navigation, search

General information

This article will cover all things Hall-C Moller. Currently, this includes all of the Qweak information. Information for past experiments can be found in the experiments section. This page is currently under construction, so please come back soon for more, or update me.

Documentation and references

The following are useful references:

Hardware related information (under construction)


The following is a plot showing a good tune of the apparatus. It was taken Jan 30, 2011 from run 700, with conditions: hallc:p is 1159 MeV, with a 1um target.

Moller plot jan30 2011.png

The preferred Moller set-up for Qweak Run II are as follows:

Quad 1: 93.5 A
Quad 3: 126 A
Col #		Position (cm)
1		-1.60
2		-1.60
3		-4.10
4		-4.10
5		Out of beam
6		-3.40
7		-3.20


For Qweak Run I, the set-up was:

Quad 1: 93 A
Quad 3: 129 A
Col #		Position (cm)
1		-1.55
2		-1.55
3		-4.05
4		-4.15
5		Out of beam
6		-3.20
7		-3.20

Hardware configuration & changes

May 10, 2012 We left the HV on during a beam position scan. We think the beam accidentally scraped somewhere, spraying the detectors. The right detector tripped off ,but the left detector didn't. It appears that the gain was screwed up, so we upped the HV. As the study continued it appeared to return to normal. The right is unchanged.

New HV Old HV
Left 1720 1640
Right --- 1685




Nov 30, 2011 Increased most of the scintillator voltages during our commissioning gain matching.

New HV Old HV
Left 1640 1630
Right 1685 1655



NB: Up until last run in December (date tbd)we used BCM 2 with gain from Kat's Faraday cup measurements.

  • (Jan 30, 2011) Increased the left and right detector high voltages, as seen below. This is a significant change in the voltages since tune-up, and is evidence of pmt radiation effects.
New HV Old HV
Left 1630 1480
Right 1655 1520



  • We've been having trouble with large charge asymmetries after switching the IHWP. To fix this, we've been slightly altering the Pockel cell HV. Apparently the wiki didn't save the numbers when I put them in earlier, so I'll put the ones I don't have with me in tomorrow.
IHWP IN OUT
POS 43886 42886
NEG 40291 41291



Experiments